# Generative Modelling of Structurally Constrained Graphs

### v1

@@ Appendix C Incremental Algorithms
The best performing full property satisfaction algorithm known for planarity is $O(n)$ [29], while its fastest known incremental test has amortized running time of $O(\alpha(q,n))$ [42] (“almost constant” complexity), where $q$ is the total number of operations (edge queries and insertions), and $\alpha$ denotes the inverse-Ackermann function.
@@
In generic undirected graphs (our case), the usual full tests via DFS/BFS have a complexity of $O(|X|+|E|)$, i.e., the algorithms have to traverse the full graph to reject the existence of any cycle. However, for the dynamic case, given that $G^{t}$ has no cycles, we can only check if the added edges, $E_{\text{added}}$, connect nodes already in the same connected component. This check can be efficiently performed if we keep an updated hashtable that maps each node to the index of the connected component it belongs at that iteration (an isolated node is a connected component) and another one with all the nodes belonging to each connected component. Whenever there is a new edge proposed, we check if the nodes are already in the same connected component. If not, we insert the edge and update the two hashtables accordingly; otherwise, we reject the edge since it would create a cycle. Therefore, the cycle check can be done in $O(|E_{\text{added}}|)$.
@@
The optimal full property satisfaction algorithm has a complexity of $O(|X|)$ since it has to perform a degree check across all nodes. The incremental version is naturally just a quick check for nodes that are vertices of $E_{\text{added}}$. Again, if we keep an updated hashtable with the degree of each node, this can be quickly performed in $O(E_{\text{added}})$.
@@ D.1 Training Details
As mentioned in Section 4, we follow the splits originally proposed for each of the unattributed datasets (lobster [45], planar [53], and tree [6]): 80% of the graphs are used in the training set and the remaining 20% are allocated to the test set. We use 20% of the train set as validation set. We note that for the lobster dataset, the original splits provided in the open-source code from  Liao et al. [45] use the validation set as a subset of the train set, i.e., all the samples in the validation set are used to train. In contrast, we follow Martinkus et al. [53]’s protocol, isolating completely the validation samples from the train (again, 20% of the train split). In any case, the test splits are coincident between our approach and the one from Liao et al. [45]. For the digital pathology datasets, we follow the same protocol.
@@ D.2 Resources
All our experiments were run in individual Nvidia V100 32Gb GPUs. We present the training and sampling times (total of 5 runs, 100 generated graphs each) for each dataset for the diffusion models in Table 3.
@@ D.3 Runtimes
A major advantage of our framework is that it does not interfere with the training of the diffusion model, preserving its efficiency, as discussed in the first paragraph of Section 3.5. Therefore, there is no overburden in the training time caused by ConStruct. For this reason, in this section we only analyse the different sampling runtimes. In particular, we track the sampling times of DiGress+ and the ones of ConStruct with and without the efficiency boosting components described in Section 3.5 (edge blocking hashtable and incremental property satisfaction algorithm).
Additionally, a natural procedure to ensure 100% constraint verification solely using DiGress+ is to first directly perform unconstrained generation and then applying a validation process to filter out the ones that do not verify the constraint. This a posteriori filtering requires a full graph property check, run only once after the graph has been generated but that is thus more computationally expensive when compared to their incremental versions employed by ConStruct. The complexity comparison between a full graph property check ran only once vs an incremental check for each added edge is property specific. However, the main bottleneck of the a posteriori filtering is that it wastes computational resources in case of graph rejection (i.e., in the case of a graph not verifying the property, all the resources used in its generation and full graph property checking are wasted) and requires restarting the sampling from zero again with an additional property check at the end. Furthermore, this procedure has to be performed sequentially, since we can only check the graphs constraint satisfaction after their generation. ConStruct avoids such redundancy and, thus, waste of computation by generating property satisfying graphs by design: throughout the reverse process, we know that that the previous graph verified the property, so we can just check property satisfaction for the newly added edges (via incremental algorithm) that have not been checked yet (via edge blocking hashtable).
@@
Table 4: Runtimes comparison. We present the results for the methods we ran over five sampling runs in the format mean ± standard error of the mean. For each run, we generated 100 graphs. All our experiments were run in individual Nvidia V100 32Gb GPUs. ConStruct [efficient] uses the edge blocking hashtable and the incremental version of the property satisfaction algorithm, while ConStruct [baseline] does not. DiGress+ refers to regular unconstrained generation, while DiGress+ [rejection] applies a posteriori filtering of unconstrained generation until we get the intended amount of graphs.
@@
ConStruct also outperforms the a posteriori filtering of unconstrained generation. In this case, we used an unconstrained model that generates constraint satisfying properties 97% (DiGress+ in Table 1 of the paper), therefore largely benefitting the unconstrained model. For example, if we considered the digital pathology setting, where we can have only 6.6% (see Table 2 of paper, high TLS dataset, DiGress+) of the generated graphs with the unconstrained model satisfying the constraint, the amount of wasted computation would be dramatically larger, implying a much worse runtime. In such setting, ConStruct is approximately 12 times more efficient in generating valid graphs than DiGress+. Additionally, these gaps in efficiency may become particularly critical in settings where the amount of generated graphs is much larger, as is the case for molecular generation (2 orders of magnitude larger, see Appendix G).
@@ E.1 Statistics
Table 5: Synthetic dataset statistics. ##\##Train, ##\##Val and ##\##Test denote the number of graphs considered in the train, validation and test splits, respectively
@@ E.2 Compared Methods
  • two spectrally conditioned methods: SPECTRE [53] is a GAN-based approach and HSpectre [6] consists of an iterative local expansion method that takes advantage of a score-based formulation;
@@
  • EDGE [9] is a degree-guided scalable discrete diffusion method (more details in Section 2);
@@ F.1 Digital Pathology Primer
Digital pathology consists of an advanced form of pathology that involves digitizing tissue slides into whole-slide images (WSI), allowing for computer-based analysis and storage. Deep learning approaches quickly integrated digital pathology processing methods, primarily focusing on extracting image-level representations for tasks such as slide segmentation and structure detection. These have also been used for downstream tasks such as cancer grading, or survival prediction [5, 69]. However, existing image-based approaches face challenges with the size WSIs, requiring their patching. This procedure raises a trade-off between the context and the patch size provided to the model. Moreover, image-based deep learning lacks efficient representations of biological entities and their relations, resulting in less interpretable models. Recently, entity-graph based approaches have emerged as a promising alternative to evade such limitations [34, 2]. These graphs are built by directly assigning nodes to biological entities and modelling their interactions with edges [25, 34], providing enhanced predictive performance and interpretability [35, 80].
Importantly, most of the deep learning contributions in the digital pathology realm have been in the discriminative setting. However, digital pathology could profoundly benefit from the development of generative formulations in several dimensions: first, there is is a lack of high-quality annotated samples, mostly due to their heavy ethical and privacy regulation. Besides, collecting these samples is remarkably costly both economically and in terms of time and labor required [34]. Most of the discriminative approaches are also instance-based. The developed models then become highly sensitive to distribution shifts, which is a common challenge across biomedical datasets, for instance due to batch effects [20]. The development of generative models in digital pathology can address these limitations by enabling both the generation of synthetic data and distribution-based characterisations of the data. Even though some approaches have been carried out using image-based methods (e.g., GANs [38] or even diffusion models [55]), these lack the advantages of graph-based approaches. To the best of our knowledge, graph-based generative modelling in digital pathology has only been explored by Madeira et al. [51]. Despite the promising results for data augmentation settings, only an off-the-shelf graph generative model (DiGress) is explored and in a proprietary dataset.
@@ F.2 Building Whole-slide Cell Graphs
Regarding edges, we followed the typical procedure for cell-graphs in digital pathology [34, 35, 2, 80]: first we used Delaunay triangulation on the cell positions to build them. Then, we discard edges longer than 25 $\mu m$. We note that we obtain different graphs than the ones considered by Danenberg et al. [14]. In terms of dimensionality, we obtain graphs with $b=9$ and $c=1$. We focus on the generation of simple yet biologically meaningful structures, Tertiary Lymphoid Structures (TLSs), further described in Section F.3. Thus, we extract 4-hop non-overlapping subgraphs centered at nodes whose class is “B-cell" from the whole-slide graphs.
@@ F.3 Tertiary Lymphoid Structures
Tertiary Lymphoid Structures (TLSs) are simple yet biologically meaningful structures. Structurally, TLSs are well-organized biological entities where clusters of B-cells are enveloped by supporting T-cells. Typically observed in ectopic locations associated with chronic inflammation [60, 67], these structures have been linked to extended disease-free survival in cancer [27, 43, 17, 57, 67], thus constituting an important indicator for medical prognosis in cancer. Since these are small structures when compared with the size of whole-slide graph, we extract non-overlapping 4-hop subgraphs centered at nodes whose class is “B-cell” from the WSI graphs. This procedure is illustrated in  Figure 4.
@@
As mentioned in Section 4.2, the TLS content of a cell graph can be quantified using the TLS embedding, $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$ [67, 51]. This TLS-like organization metric considers only edges between B and T-cells and classifies them into several categories: $\alpha$ edges link two cells of the same type, while $\gamma_{j}$ edges connect a B to a T-cell, where $j$ is the number of B-cell neighbors of the B-cell vertex (see Figure 4). Therefore, the entry $i$ of $\kappa$ is defined as the proportion of its $\gamma$ edges whose index is larger than $i$:
@@
where $|E_{\text{BT}}|$, $|E_{\alpha}|$, and $|E_{\gamma_{j}}|$ correspond to the number of edges between B and T-cells, of $\alpha$ edges and of $\gamma_{j}$ edges in a given graph, $G$. By definition, note that the entries of $\kappa$ take values between 0 and 1 and are monotonically non-increasing with $i$.
@@ F.4 Statistics of digital pathology datasets
Table 7: Digital pathology datasets statistics. ##\##Train, ##\##Val and ##\##Test denote the number of graphs considered in the train, validation and test splits, respectively.
@@
Figure 5: Distributions of the TLS embedding entries for high TLS (left) and the low TLS content (right) datasets.
@@ F.5 Baseline Method for Digital Pathology
The non deep learning method used as baseline for the digital pathology dataset follows Madeira et al. [51]. This model learns three distributions by counting the proportions of given events in the train dataset. In particular:
@@
To sample a new graph, we first sample a number of nodes for the graph from the first distribution. Then, for each of those nodes sample a cell phenotype from the second distribution. Finally, between every pair of cells, we sample an edge type given the two phenotypes previously sampled. This sampling algorithm is described in Algorithm 4.
@@ G.1 Exploring Planarity
  • FCD - Fréchet ChemNet Distance [62], similar to Fréchet Inception Distance (FID) but for molecules, represented as SMILES. This metric evaluates sample quality in unconstrained settings.
@@
  • Valid - proportion of valid molecules (checked using valencies). This metric evaluates sample validity.
In particular, we explore planarity as target structural property in the molecular setting, as previously suggested by recent works for the discriminative tasks [18]. Therefore, we also include the proportion of planar molecules from the generated set as an evaluation metric. As a remark, QM9 and MOSES are exclusively composed of planar molecules. GuacaMol contains 3 non-planar molecules out of 1273104 molecules in the train set and 3 non-planar molecules out of 238706 molecules in the test set. We considered these non-planar examples negligible in model training/evaluation, thus fully preserving the original dataset. The validation set has 79568 planar molecules. The results are shown in Table 9.
Table 9: Graph diffusion performance on molecular generation. The constraining property used for ConStruct is planarity. We present the results for the methods we ran over five sampling runs in the format mean ± standard error of the mean. For each run, we generated 10000, 25000, and 18000 generated molecules for QM9, MOSES, and GuacaMol, respectively, following the protocol from Vignac et al. [77].
@@
We only observe an incremental improvement in the planarity satisfaction of the output graphs, since DiGress+ learns to almost always generate planar graphs. As a consequence, the constrained model ends up being marginally less expressive than the unconstrained model, as extensively discussed in Section H.1. This impacts both sample quality and sample validity. In fact, while for the discriminative setting, the main aim is to design fully expressive architectures for classes of graphs as broad as possible, as demonstrated by PlanE [18] for planar graphs, for constrained generation, this is not the case. As exemplified in this case, planarity is too loose of a constraint, since the atoms composing molecules typically have low degrees and it becomes highly unlikely for the unconstrained diffusion model to violate planarity. This ends up slightly harming the performance of the constrained generative model without bringing the benefits of increased validity, as observed in Section 4.
As a side note for the interested reader, the low values of novelty for QM9 are a result of the nature of this dataset, which consists of an exhaustive enumeration of small molecules satisfying a given set of properties [76, 77]. Therefore, there is small room for the generation of new molecules within such domain.

### v2

@@ Appendix C Incremental Algorithms
The best performing full property satisfaction algorithm known for planarity is $O(n)$ [30], while its fastest known incremental test has amortized running time of $O(\alpha(\mathfrak{q},n))$ [43] (“almost constant” complexity), where $\mathfrak{q}$ is the total number of operations (edge queries and insertions), and $\alpha$ denotes the inverse-Ackermann function.
@@
In generic undirected graphs (our case), the usual full tests via DFS/BFS have a complexity of $O(n+|E|)$, i.e., the algorithms have to traverse the full graph to reject the existence of any cycle. However, for the dynamic case, given that $G^{t}$ has no cycles, we can only check if the added edges, $E_{\text{added}}$, connect nodes already in the same connected component. This check can be efficiently performed if we keep an updated hashtable that maps each node to the index of the connected component it belongs at that iteration (an isolated node is a connected component) and another one with all the nodes belonging to each connected component. Whenever there is a new edge proposed, we check if the nodes are already in the same connected component. If not, we insert the edge and update the two hashtables accordingly; otherwise, we reject the edge since it would create a cycle. Therefore, the cycle check can be done in $O(|E_{\text{added}}|)$.
@@
The optimal full property satisfaction algorithm has a complexity of $O(n)$ since it has to perform a degree check across all nodes. The incremental version is naturally just a quick check for nodes that are vertices of $E_{\text{added}}$. Again, if we keep an updated hashtable with the degree of each node, this can be quickly performed in $O(|E_{\text{added}}|)$.
@@ D.1 Training Details
As mentioned in Section 4, we follow the splits originally proposed for each of the unattributed datasets (lobster [46], planar [54], and tree [7]): 80% of the graphs are used in the train set and the remaining 20% are allocated to the test set. We use 20% of the train set as validation set. We note that for the lobster dataset, the original splits provided in the open-source code from Liao et al. [46] use the validation set as a subset of the train set, i.e., all the samples in the validation set are used to train. In contrast, we follow Martinkus et al. [54]’s protocol, isolating completely the validation samples from the train (again, 20% of the train split). In any case, the test splits are coincident between our approach and the one from Liao et al. [46]. For the digital pathology datasets, we follow the same protocol.
@@ D.2 Resources
All our experiments were run in a single Nvidia V100 32Gb GPUs. We present the training times of the diffusion model for each dataset in Table 3.
@@ D.3 Runtimes
A major advantage of our framework is that it does not interfere with the training of the diffusion model, preserving its efficiency. Therefore, there is no overburden in the training time caused by ConStruct. For this reason, in this section we only analyse the different sampling runtimes. In particular, we track the sampling times of DiGress+ and the ones of ConStruct with and without the efficiency boosting components described in Section 3.5 (edge blocking hashtable and incremental property satisfaction algorithm).
Additionally, a natural procedure to ensure 100% constraint verification solely using DiGress+ is to first directly perform unconstrained generation and then applying a validation process to filter out the ones that do not verify the constraint. This a posteriori filtering requires a full graph property check, run only once after the graph has been generated but that is thus more computationally expensive when compared to their incremental versions employed by ConStruct. The complexity comparison between a full graph property check ran only once vs an incremental check for each added edge is property specific. However, the main bottleneck of the a posteriori filtering is that it wastes computational resources in case of graph rejection (i.e., in the case of a graph not verifying the property, all the resources used in its generation and full graph property checking are wasted) and requires restarting the sampling from zero again with an additional property check at the end. Furthermore, this procedure has to be performed sequentially, since we can only check the graphs constraint satisfaction after their generation. ConStruct avoids such redundancy and, thus, waste of computation by generating property satisfying graphs by design: throughout the reverse process, we know that the previous graph verified the property, so we can just check property satisfaction for the newly added edges (via incremental algorithm) that have not been checked yet (via edge blocking hashtable).
@@
Table 4: Runtimes comparison. We performed five sampling runs for each method and present their results in the format mean ± standard error of the mean. For each run, we generated 100 graphs. All our experiments were run in a single Nvidia V100 32Gb GPUs. ConStruct [efficient] uses the edge blocking hashtable and the incremental version of the property satisfaction algorithm, while ConStruct [baseline] does not. DiGress+ refers to regular unconstrained generation, while DiGress+ [rejection] applies a posteriori filtering of unconstrained generation until we get the intended amount of graphs.
@@
ConStruct also outperforms the a posteriori filtering of unconstrained generation. In this case, we used an unconstrained model that generates constraint satisfying properties 97% (DiGress+ in Table 1 of the paper), therefore largely benefitting the unconstrained model. For example, if we considered the digital pathology setting, where we can have only 6.6% (see Table 2, high TLS dataset, DiGress+) of the generated graphs with the unconstrained model satisfying the constraint, the amount of wasted computation would be dramatically larger, implying a much worse runtime. In such setting, ConStruct would be approximately 12 times more efficient in generating valid graphs than DiGress+. Additionally, this gap in sampling efficiency may become particularly critical in settings where the amount of generated graphs is much larger, as is the case for molecular generation (two orders of magnitude greater than in the settings with synthetic datasets, see Appendix G).
@@ E.1 Statistics
Table 5: Synthetic dataset statistics. ##\##Train, ##\##Val and ##\##Test denote the number of graphs considered in the train, validation and test splits, respectively.
@@ E.2 Compared Methods
  • two spectrally conditioned methods: SPECTRE [54] is a GAN-based approach and HSpectre [7] consists of an iterative local expansion method that takes advantage of a score-based formulation for intermediate steps;
@@
  • EDGE [10] is a degree-guided scalable discrete diffusion method (more details in Section 2).
@@ F.1 Digital Pathology Primer
Digital pathology consists of an advanced form of pathology that involves digitizing tissue slides into whole-slide images (WSI), allowing for computer-based analysis and storage. Deep learning approaches quickly integrated digital pathology processing methods, primarily focusing on extracting image-level representations for tasks such as slide segmentation and structure detection. These have also been used for downstream tasks such as cancer grading, or survival prediction [6, 71]. However, existing image-based approaches face challenges with the sizes of WSIs, requiring their patching. This procedure raises a trade-off between the context and the size of the patch provided to the model. Moreover, image-based deep learning lacks efficient representations of biological entities and their relations, resulting in less interpretable models. Recently, entity-graph based approaches have emerged as a promising alternative to evade such limitations [36, 2]. These graphs are built by directly assigning nodes to biological entities and modelling their interactions with edges [26, 36], providing enhanced predictive performance and interpretability [35, 82].
Importantly, most of the deep learning contributions in the digital pathology realm have been in the discriminative setting. However, digital pathology could profoundly benefit from the development of generative formulations in several dimensions: first, there is is a lack of high-quality annotated samples, mostly due to their heavy ethical and privacy regulation. Besides, collecting these samples is remarkably costly, both economically and in terms of time and labor required [36]. Most of the discriminative approaches are also instance-based. The developed models then become highly sensitive to distribution shifts, which is a common challenge across biomedical datasets, for instance due to batch effects [21]. The development of generative models in digital pathology can address these limitations by enabling both the generation of synthetic data and distribution-based characterisations of the data. Even though some approaches have been carried out using image-based methods (e.g., GANs [39] or even diffusion models [56]), these lack the advantages of graph-based approaches. To the best of our knowledge, graph-based generative modelling in digital pathology has only been explored by Madeira et al. [52]. Despite the promising results for data augmentation settings, only an off-the-shelf graph generative model (DiGress) is explored and in a proprietary dataset.
@@ F.2 Building Whole-slide Cell Graphs
Regarding edges, we followed the typical procedure for cell-graphs in digital pathology [36, 35, 2, 82]: first we used Delaunay triangulation on the cell positions to build them. Then, we discard edges longer than 25 $\mu m$. We note that we obtain different graphs than the ones considered by Danenberg et al. [15]. In terms of dimensionality, we obtain graphs with $b=9$ and $c=1$. We focus on the generation of simple yet biologically meaningful structures, Tertiary Lymphoid Structures (TLSs), further described in the next section. Thus, we extract 4-hop non-overlapping subgraphs centered at nodes whose class is “B-cell" from the whole-slide graphs.
@@ F.3 Tertiary Lymphoid Structures
Tertiary Lymphoid Structures (TLSs) are simple yet biologically meaningful structures. Structurally, TLSs are well-organized biological entities where clusters of B-cells are enveloped by supporting T-cells. Typically observed in ectopic locations associated with chronic inflammation [61, 69], these structures have been linked to extended disease-free survival in cancer [28, 44, 18, 58, 69], thus constituting an important indicator for medical prognosis in cancer. Since these are small structures when compared with the size of whole-slide graph, we extract non-overlapping 4-hop subgraphs centered at nodes whose class is “B”, corresponding to B cells, from the WSI graphs. This procedure is illustrated in Figure 4.
@@
As mentioned in Section 4.2, the TLS content of a cell graph can be quantified using the TLS embedding, $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$ [69, 52]. This TLS-like organization metric considers only edges between B and T-cells and classifies them into several categories: $\alpha$ edges link two cells of the same type, while $\gamma_{j}$ edges connect a B to a T-cell, where $j$ is the number of B-cell neighboring the B-cell vertex (see Figure 4). Therefore, the entry $i$ of $\kappa$ is defined as the proportion of its $\gamma$ edges whose index is larger than $i$:
@@
where $|E_{\text{BT}}|$, $|E_{\alpha}|$, and $|E_{\gamma_{j}}|$ correspond to the number of edges whose both vertices are B or T-cells, of $\alpha$ edges and of $\gamma_{j}$ edges in a given graph, $G$. Note that, by definition, the entries of $\kappa$ take values between 0 and 1 and are monotonically non-increasing with $i$.
@@ F.4 Statistics of digital pathology datasets
Table 7: Digital pathology datasets statistics. Here, we report the same stats as in Table 5. ##\##Train, ##\##Val and ##\##Test denote the number of graphs considered in the train, validation, and test splits, respectively.
@@
Figure 5: Distributions of the TLS embedding entries for the low TLS (left) and the high TLS (right) datasets.
@@ F.5 Baseline Method for Digital Pathology
The non-deep learning method used as baseline for the digital pathology dataset follows Madeira et al. [52]. This model learns three distributions by counting the frequencies of given events in the train dataset. In particular:
@@
To sample a new graph, we first sample a number of nodes for the graph from the first distribution. Then, for each of those nodes sample a cell phenotype from the second distribution. Finally, between every pair of cells, we sample an edge type given the two phenotypes previously sampled from the third distribution. This sampling algorithm is described in Algorithm 4.
@@ G.1 Exploring Planarity
  • FCD - Fréchet ChemNet Distance [63], similar to Fréchet Inception Distance (FID) but for molecules, represented as SMILES. This metric evaluates the similarity between the generated and test molecule sets, providing an indicator for the sample quality in unconstrained settings;
@@
  • Valid - proportion of generated molecules that are valid. This metric evaluates sample validity.
In particular, we explore planarity as target structural property in the molecular setting, as previously suggested by recent works in discriminative tasks [19]. Therefore, we also include the proportion of planar molecules from the generated set as an evaluation metric. As a remark, QM9 and MOSES are exclusively composed of planar molecules. GuacaMol contains 3 non-planar molecules out of 1273104 molecules in the train set and 3 non-planar molecules out of 238706 molecules in the test set. We considered these non-planar examples negligible in model training/evaluation, thus fully preserving the original dataset. The validation set has 79568 planar molecules. The results are shown in Table 9.
Table 9: Graph diffusion performance on molecular generation. The constraining property used for ConStruct is planarity. We performed five sampling runs for each method and present their results in the format mean ± standard error of the mean. For each run, we generated 10000, 25000, and 18000 generated molecules for QM9, MOSES, and GuacaMol, respectively, following the protocol from Vignac et al. [79]. Note that for MOSES and GuacaMol, we do not report their benchmarking metrics, as we focus on an overview comparative analysis of the two methods, DiGress+ and ConStruct. The FCD is computed using the official implementation from Preuer et al. [63].
@@
We only observe an incremental improvement in the planarity satisfaction of the output graphs of ConStruct, since DiGress+ learns to almost always generate planar graphs. As a consequence, the constrained model ends up being marginally less expressive than the unconstrained one, as extensively discussed in Section H.1. This impacts both sample quality and sample validity. In fact, while for the discriminative setting, the main aim is to design fully expressive architectures for classes of graphs as broad as possible, e.g., PlanE [19] for planar graphs, in constrained generation this is not the case. As exemplified, planarity is too loose of a constraint, since the atoms composing molecules typically have low degrees and it becomes highly unlikely for the unconstrained diffusion model to violate planarity. This ends up slightly harming the performance of the constrained generative model without bringing the benefits of increased validity, as observed in Section 4.
As a side note for the interested reader, the low values of novelty for QM9 are a result of the nature of this dataset, which consists of an exhaustive enumeration of small molecules satisfying a given set of properties [78, 79]. Therefore, there is small room for the generation of new molecules within such space.